A car that travels at 6.3 km/hr for 2.0 hrs, at 5.7 for 1.0 hr and at 6.6 for half an hour. What is its average speed?


1
Expert's answer
2022-03-11T08:27:45-0500

Solution.

"v_1=6.3 km\/hr;"

"t_1=2.0hr;"

"v_2=5.7 km\/hr;"

"t_2=1.0hr;"

"v_3=6.6 km\/hr;"

"t_3=0.5 hr;"

"v_{average}=\dfrac{l_1+l_2+l_3}{t_1+t_2+t_3};"

"l_1=v_1\sdot t_1=6.3\sdot 2.0=12.6km;"

"l_2=v_2\sdot t_2=5.7\sdot1.0=5.7km;"

"l_3=v_3\sdot t_3=6.6\sdot0.5=3.3km;"

"v_{average}=\dfrac{12.6+5.7+3.3}{2.0+1.0+0.5}=6.17 km/hr;"

Answer: "v_{average}=6.17 km\/hr."